OpenStudy (anonymous):

Use U-substitution to find the real and imaginary solutions to the following equation: x-5x^1/2+4=0

OpenStudy (anonymous):

try letting u=x^1/2 what is u^2=?

OpenStudy (anonymous):

u^2=(x^1/2)^2=? try that

OpenStudy (anonymous):

i got down to (x^1/2-4) (x^1/2-1)=0

OpenStudy (anonymous):

then i am stuck

OpenStudy (anonymous):

no, try substitution of u instead of x

OpenStudy (anonymous):

yes i did and then i substituted x^1/2 back in for u

OpenStudy (anonymous):

ah ok.... u=x^1/2 u^2=(x^1/2)^2=? x correct? then from x-5x^1/2+4=0 sub u^2-5u+4=0 now use quadratic formula to solve for u, then x

OpenStudy (anonymous):

yes then i got (u-4) (u-1)

OpenStudy (anonymous):

and then substituted x^1/2 back in for u

OpenStudy (anonymous):

ok very good, now solve for x since u=x^1/2

OpenStudy (anonymous):

so what do u do when u get to (x^1/2-4) (x^1/2-1)=0

OpenStudy (anonymous):

subtract one to the other side?

OpenStudy (anonymous):

(u-4) (u-1)=0 u=4,1 u=x^1/2 4=x^1/2 squaring both sides 4^2=(x^1/2)^2 16=x and 1=x^1/2 1=x therefore x=1 and x=16 ..ans

OpenStudy (anonymous):

let me know if u hve question

OpenStudy (anonymous):

my teacher instructed me to get to here (x^1/2-4)=(x^1/2-1) and then square both sides.

OpenStudy (anonymous):

how would i do this